### v3

# Amenable covers and relative bounded cohomology

Abstract.
We establish a relative version of Gromov’s Vanishing Theorem in the presence of amenable open covers with small multiplicity, extending a result of Li, Löh and Moraschini. Our approach relies on Gromov’s theory of multicomplexes.
Key words and phrases:
amenable groups, bounded cohomology, multicomplexes
1000 Mathematics Subject Classification:
55N10, 55U10, 57N65
1. Introduction
Let $X$ be a topological space. Gromov’s Vanishing Theorem [Gro82] affirms that, if $X$ admits an amenable open cover of multiplicity $k$, then the comparison map $H^{n}_{b}(X;\mathbb{R})\rightarrow H^{n}(X;\mathbb{R})$ from bounded cohomology to singular cohomology of $X$ vanishes in every degree $n\geq k$. Several proofs of this result are available in the literature, whose techniques range between Gromov’s theory of multicomplexes [Gro82, Fri22, FM23], sheaf theory [Iva87, Iva], equivariant nerves and classifying spaces for families [LS20] and homotopy theory [Rap24]. A generalization of Gromov’s result to the relative setting has been provided by Li, Löh and Moraschini in [LLM] and by Raptis in [Rap24]. In this paper we extend the result of Li, Löh and Moraschini in two directions, following Gromov’s approach via multicomplexes. We refer the reader to Remark 1.2 for a detailed discussion on the relationship between our results and the ones in [LLM]. A pair of topological spaces $(X,A)$ is called triangulable if there exists a pair of simplicial complexes $(T,S)$ such that $(X,A)=(|T|,|S|)$.
Theorem 1.
Let $(X,A)$ be a triangulable pair and assume that the kernel of the morphism $\pi_{1}(A\hookrightarrow X,x)$ is amenable for every $x\in A$. Let $\mathcal{U}$ be an amenable cover of $X$ by path-connected open subsets such that:
  1. $(\mathrm{RC1})$
    For every $U\in\mathcal{U}$, $U\cap A$ is path-connected;
  2. $(\mathrm{RC2})$
    For every $U\in\mathcal{U}$, the inclusion
    $\mathrm{im}(\pi_{1}(U\cap A,x)\rightarrow\pi_{1}(X,x))\hookrightarrow\mathrm{im}(\pi_{1}(U,x)\rightarrow\pi_{1}(X,x))$
    is an isomorphism for every $x\in U\cap A$.
Then the comparison map $\mathrm{comp}^{n}\colon H^{n}_{b}(X,A;\mathbb{R})\rightarrow H^{n}(X,A;\mathbb{R})$ vanishes for every $n\geq\mathrm{mult}(\mathcal{U})$.
Remark 1.1.
In the statement of (RC1) we follow the convention that the empty set is path-connected. Moreover, (RC2) is empty when $\mathcal{U}$ consists of $\pi_{1}$-contractible subsets (for which the map $\pi_{1}(U,x)\rightarrow\pi_{1}(X,x)$ is trivial for every $U\in\mathcal{U}$ and every $x\in U$).
Theorem 2.
Let $(X,A)$ be a triangulable pair and assume that the kernel of the morphism $\pi_{1}(A\hookrightarrow X,x)$ is amenable for every $x\in A$. Let $\mathcal{U}$ be a locally-finite amenable cover of $X$ by path-connected open subsets satisfying $(\mathrm{RC1})$ and $(\mathrm{RC2})$.
  1. $(1)$
    If $\mathcal{U}$ is weakly-convex on $A$, then the comparison map $\mathrm{comp}^{n}$ vanishes for every $n\geq\mathrm{mult}_{A}(\mathcal{U})$.
  2. $(2)$
    If $\mathcal{U}$ is convex, then for every $n\in\mathbb{N}$ there exists a map $\Theta^{n}$ such that the following diagram commutes
    $H^{n}_{b}(X,A;\mathbb{R})$$H^{n}(X,A;\mathbb{R})$$H^{n}(N(\mathcal{U}),N(\mathcal{U}_{A});\mathbb{R})$$H^{n}(|N(\mathcal{U})|,|N(\mathcal{U}_{A})|;\mathbb{R}),$$\scriptstyle{\mathrm{comp}^{n}}$$\scriptstyle{\Theta^{n}}\scriptstyle{\cong}$$\scriptstyle{H^{n}(\nu)}$
    where $N(\mathcal{U})$ is the nerve of $\mathcal{U}$, $N(\mathcal{U}_{A})$ is the nerve of the cover of $A$ induced by $\mathcal{U}$ and $\nu\colon(X,A)\rightarrow(|N(\mathcal{U})|,|N(\mathcal{U}_{A})|)$ is a nerve map.
We refer the reader to Subsection 2.1 for the definitions of $\mathrm{mult}_{A}(\mathcal{U})$, weak-convexity and convexity. Since $\dim(N(\mathcal{U}),N(\mathcal{U}_{A}))=\mathrm{mult}_{A}(\mathcal{U})-1$ [LLM, Definition 4.4], assuming convexity, we have that (2) implies (1).
Remark 1.2.
Theorem 1 and Theorem 2 are natural extensions of the Relative Vanishing Theorem by Li, Löh and Moraschini [LLM, Theorem 1.1]. Their result in fact deduces the same conclusions (1) and (2) of our Theorem 2, under the assumptions that $(X,A)$ is a CW-pair such that $A$ is $\pi_{1}$-injective in $X$ and $\mathcal{U}$ is a (not necessarily locally-finite) boundedly-acyclic open cover by path-connected subsets satisfying (RC1) and (RC2). Of course, triangulable pairs are in particular CW-pairs. However, not every CW-pair can be triangulated [FP90, Corollary 4.6.12]. Moreover, amenable open covers are in particular boundedly-acyclic, and it is not clear whether our results can be extended to this case. On the other hand, our assumption which requires the kernel of the map $\pi_{1}(A\hookrightarrow X)$ to be amenable is clearly more flexible than $\pi_{1}$-injectivity. We also underline the fact that, with respect to [LLM], we need to assume $\mathcal{U}$ to be locally finite in Theorem 2.
Via a standard duality argument [Fri17, Section 7.5], Theorem 1 and Theorem 2 imply in a straightforward way vanishing results for relative simplicial volume (see for example [LLM, Corollary 6.15]). As already noticed in [LLM, Section 6.4], these results are however strictly weaker than the main available result for relative simplicial volume [LMR22, Theorem 3.13], which is based on Gromov’s vanishing theorem for non-compact manifolds [FM23, Corollary 11]. We conclude with a discussion on the optimality of our assumptions.
Remark 1.3.
Let $S$ be an oriented compact connected surface of genus one with one boundary component. It is well known that the relative simplicial volume $\|S,\partial S\|$ of $S$ is positive [Gro82], hence the comparison map $\mathrm{comp}^{2}\colon H^{2}_{b}(S,\partial S;\mathbb{R})\rightarrow H^{2}(S,\partial S;\mathbb{R})$ is non-zero [Fri17, Section 7.5]. Moreover, the boundary $\partial S$ is $\pi_{1}$-injective in $S$. Therefore, the open covers of $S$ in Figure 1 show that the regularity assumptions (RC1) and (RC2) of Theorem 1 are optimal. We refer to [LLM, Remark 6.16] for a discussion on the optimality of the assumptions of Theorem 2.
Figure 1. These open covers of $S$ are amenable and have multiplicity 2. The one on the left satisfies (RC1), but not (RC2), the one on the right satisfies (RC2), but not (RC1).
Acknowledgments
We wish to thank Marco Moraschini for suggesting the topic of this paper. We are grateful to Roberto Frigerio for his guidance. We would also like to thank Federica Bertolotti, Filippo Bianchi, Kevin Li and Francesco Milizia for useful discussions.
2. Preliminaries
2.1. Open covers
Let $X$ be a topological space. An open cover $\mathcal{U}$ of $X$ is a set of open subsets of $X$ such that $\bigcup_{U\in\mathcal{U}}U=X$. An open cover $\mathcal{U}$ of $X$ is called amenable if, for every $U\in\mathcal{U}$ and every $x\in U$, we have that $\mathrm{im}(\pi_{1}(U\hookrightarrow X,x))$ is an amenable subgroup of $\pi_{1}(X,x)$. The multiplicity of $\mathcal{U}$ is the supremum of the set of natural numbers $k\in\mathbb{N}_{\geq 1}$ such that $U_{1}\cap\dots\cap U_{k}\neq\emptyset$ for some pairwise distinct $U_{1},\dots,U_{k}\in\mathcal{U}$. We say that $\mathcal{U}$ is convex if, for every $k\in\mathbb{N}$ and every $U_{1},\dots,U_{k}\in\mathcal{U}$, the intersection $U_{1}\cap\dots\cap U_{k}$ is path-connected. The nerve $N(\mathcal{U})$ of $\mathcal{U}$ is the simplicial complex whose set of vertices is the set $\mathcal{U}$ itself and pairwise-different $U_{0},\dots,U_{n}\in\mathcal{U}$ span an $n$-simplex in $N(\mathcal{U})$ if $U_{0}\cap\dots\cap U_{n}\neq\emptyset$. Of course, $\dim(N(\mathcal{U}))=\mathrm{mult}(\mathcal{U})-1$.
Let $(X,A)$ be a pair of topological spaces. The following notions were introduced by Li, Löh and Moraschini in [LLM]. An open cover $\mathcal{U}$ of $X$ is called weakly-convex on $A$ if, for every $k\in\mathbb{N}$ and every $U_{1},\dots,U_{k}\in\mathcal{U}$ with $U_{1}\cap\dots\cap U_{k}\cap A\neq\emptyset$, each path-connected component of $U_{1}\cap\dots\cap U_{k}$ intersects $A$. Convex open covers are also weakly convex. The relative multiplicity $\mathrm{mult}_{A}(\mathcal{U})$ of $\mathcal{U}$ (with respect to $A$) is defined as the supremum of the set of natural numbers $k\in\mathbb{N}_{\geq 1}$ such that $U_{1}\cap\dots\cap U_{k}\neq\emptyset$ and $U_{1}\cap\dots\cap U_{k}\cap A=\emptyset$ for some pairwise distinct $U_{1},\dots,U_{k}\in\mathcal{U}$. Of course, $\mathrm{mult}_{A}(\mathcal{U})$ is a lower bound of $\mathrm{mult}(\mathcal{U})$.
The regularity conditions (RC1) and (RC2) are used in our proof of Theorem 1 via the following lemma.
Lemma 2.1.
Let $(X,A)$ be a pair of topological spaces and let $\mathcal{U}$ be an cover of $X$ by path-connected open subsets. Then $\mathcal{U}$ satisfies the following conditions:
  1. $(\mathrm{RC1})$
    For every $U\in\mathcal{U}$, $U\cap A$ is path-connected;
  2. $(\mathrm{RC2})$
    For every $U\in\mathcal{U}$, the inclusion
    $\mathrm{im}(\pi_{1}(U\cap A,x)\rightarrow\pi_{1}(X,x))\hookrightarrow\mathrm{im}(\pi_{1}(U,x)\rightarrow\pi_{1}(X,x))$
    is an isomorphism for every $x\in U\cap A$;
if and only if, for every $U\in\mathcal{U}$ and for every path $\gamma\colon[0,1]\rightarrow U$ with endpoints in $U\cap A$, there exists a path $\lambda\colon[0,1]\rightarrow U\cap A$ such that $\lambda$ is homotopic in $X$ to $\gamma$ relative to the endpoints.
Proof.
Assume that $\mathcal{U}$ satisfies (RC1) and (RC2). Given $U\in\mathcal{U}$ and $\gamma$ as above, by (RC1) there exists a path $\varepsilon\colon[0,1]\rightarrow U\cap A$ from $\gamma(0)$ to $\gamma(1)$. The homotopy class of the loop $\gamma*\bar{\varepsilon}$ identifies an element of $\mathrm{im}(\pi_{1}(U\rightarrow X,\gamma(0)))$, hence, by (RC2), there exists a loop $\delta\colon[0,1]\rightarrow U\cap A$ which is homotopic in $X$ to $\gamma*\bar{\varepsilon}$ relative to the endpoints. Therefore, we can set $\lambda\coloneqq\delta*\varepsilon$. The converse implication is straightforward. ∎
2.2. Relative bounded cohomology of pairs
Let $(X,A)$ be a pair of topological spaces. Let $C^{\bullet}(X)$ and $C^{\bullet}(A)$ denote the real singular cochain complexes of $X$ and $A$ respectively. We denote by $C^{\bullet}(X,A)$ the kernel of the restriction map $C^{\bullet}(X)\rightarrow C^{\bullet}(A)$. The real bounded cohomology $H^{\bullet}_{b}(X,A)$ of $(X,A)$ is the cohomology of the subcomplex $C^{\bullet}_{b}(X,A)\subseteq C^{\bullet}(X,A)$ given by bounded cochains, where a cochain $f\in C^{n}(X)$ is called bounded if
$\lVert f\rVert_{\infty}=\sup\bigl\{|f(\sigma)|,\;\sigma\mbox{ is a singular $n$-simplex}\bigr\}<\infty.$
The inclusion of complexes $C^{\bullet}_{b}(X,A)\subseteq C^{\bullet}(X,A)$ induces the so-called comparison map $\mathrm{comp}^{n}\colon H^{n}_{b}(X,A)\rightarrow H^{n}(X,A)$.
2.3. Multicomplexes
Multicomplexes are simplicial structures introduced by Gromov in [Gro82]. A multicomplex can be defined as a $\Delta$-complex that is both regular and unordered (see [Hat]) i.e. a $\Delta$-complex whose simplices have unordered and distinct vertices, where multiple simplices may share the same set of vertices. Every simplicial complex is of course also a multicomplex. We refer the reader to [FM23] for a more precise discussion of the notion of multicomplex.
A pair of multicomplexes is a pair $(K,L)$, where $L$ is a submulticomplex of $K$. A simplicial map between pairs of multicomplexes $f\colon(K,L)\rightarrow(K^{\prime},L^{\prime})$ is given by a simplicial map $f\colon K\rightarrow K^{\prime}$ such that $f(L)\subseteq L^{\prime}$. We denote by $\textup{Aut}(K,L)$ the group of simplicial automorphisms of the pair $(K,L)$. Given a group $G$, a simplicial group action $G\curvearrowright(K,L)$ is given by a group homomorphism $G\rightarrow\textup{Aut}(K,L)$.
Let $K$ be a multicomplex. An algebraic $n$-simplex of $K$ is a pair
$\sigma=(\Delta,(v_{0},\dots,v_{n})),$
where $\Delta$ is a $k$-simplex of $K$, and the set $\{v_{0},\dots,v_{n}\}$ is the set of vertices of $\Delta$. We do not require the elements of the ordered $(n+1)$-tuple $(v_{0},\dots,v_{n})$ to be pairwise distinct. However, since $\Delta$ has exactly $k+1$ vertices, one has $k\leq n$. We denote by $C_{n}(K)$ the vector space generated by algebraic $n$-simplices. For every $i\in\{0,\dots,n\}$, the $i$-th face $\partial_{n}^{i}\sigma$ of $\sigma$ is given by
$\partial_{n}^{i}\sigma=(\Delta^{\prime},(v_{0},\dots,\widehat{v}_{i},\dots,v_{n}))\in C_{n-1}(K),$
where $\Delta^{\prime}=\Delta$, if $\{v_{0},\dots,v_{n}\}=\{v_{0},\dots,\widehat{v_{i}},\dots,v_{n}\}$, and $\Delta^{\prime}$ is the unique face of $\Delta$ with vertices $\{v_{0},\dots,\widehat{v_{i}},\dots,v_{n}\}$, otherwise. The boundary operator $\partial_{n}\colon C_{n}(K)\rightarrow C_{n-1}(K)$, defined as the alternating sum of the faces $\partial_{n}^{i}$, endows $C_{\bullet}(K)$ with the structure of a chain complex. Let $C^{\bullet}(K)$ (resp. $C^{\bullet}_{b}(K)$) be the complex of (bounded) real simplicial cochains on $K$, and let $H^{\bullet}(K)$ (resp. $H^{\bullet}_{b}(K)$) be the corresponding cohomology module.
If $L$ is a submulticomplex of $K$, the inclusion $C_{\bullet}(L)\subseteq C_{\bullet}(K)$ induces a short exact sequence of chain complexes
$0\rightarrow C^{\bullet}_{b}(K,L)\rightarrow C^{\bullet}_{b}(K)\rightarrow C^{\bullet}_{b}(L)\rightarrow 0.$
We denote by $H^{\bullet}_{b}(K,L)$ the cohomology of the complex $C^{\bullet}_{b}(K,L)$, which in turn fits into the long exact sequence
$\dots\rightarrow H^{n-1}_{b}(L)\rightarrow H^{n}_{b}(K,L)\rightarrow H^{n}_{b}(K)\rightarrow H^{n}_{b}(L)\rightarrow\dots$
A cochain $f\in C^{n}(K)$ is called alternating if, for every algebraic simplex $(\Delta,(v_{0},\dots,v_{n}))$ and every permutation $\tau$ of the set $\{0,\dots,n\}$, we have that $f(\Delta,(v_{0},\dots,v_{n}))=\text{sign}(\tau)\cdot f(\Delta,(v_{\tau(0)},\dots,v_{\tau(n)}))$. We denote by $C^{\bullet}_{b}(K)_{\textup{alt}}$ the subcomplex of alternating cochains. Since the inclusion $C^{\bullet}_{b}(K)_{\textup{alt}}\subseteq C^{\bullet}_{b}(K)$ induces an isomorphisms in cohomology [FM23, Theorem 1.9], it follows from the Five Lemma that the inclusion
$C^{\bullet}_{b}(K,L)_{\textup{alt}}=\ker(C^{\bullet}_{b}(K)_{\textup{alt}}\rightarrow C^{\bullet}_{b}(L)_{\textup{alt}})\rightarrow C^{\bullet}_{b}(K,L)$
induces an isomorphisms in cohomology. In particular, every cohomology class in $H^{\bullet}_{b}(K,L)$ can be represented by an alternating cocycle.
2.4. Invariant chains
If a group $G$ acts simplicially on a multicomplex $K$, then it induces a linear action $G\curvearrowright C_{\bullet}(K)$ on the space of algebraic simplices. This in turns induces a linear action $G\curvearrowright C^{\bullet}_{b}(K)$ on the space of bounded cochains. We denote by $C^{\bullet}_{b}(K)^{G}$ the space of $G$-invariant bounded cochains. Let $L$ be a submulticomplex of $K$ and let $H$ be a subgroup of $G$ whose action on $K$ preserves $L$. Then there is an obvious restriction map
$r_{G,H}^{\bullet}\colon C^{\bullet}_{b}(K)^{G}\rightarrow C^{\bullet}_{b}(L)^{H},$
whose kernel is $C^{\bullet}_{b}(K,L)^{G}\coloneqq C^{\bullet}_{b}(K)^{G}\cap C^{\bullet}_{b}(K,L)$. The following regularity condition for the action $G\curvearrowright K$ on $L$ was introduced in [Cap, Definition 9.1]. The action $G\curvearrowright K$ has orbits in $L$ induced by $H$ if, for every algebraic simplex $\sigma$ of $L$ and every $g\in G$ such that $g\cdot\sigma$ is an algebraic simplex of $L$, there exists $h\in H$ such that $h\cdot\sigma=g\cdot\sigma$.
[…]
Lemma 2.2.
Let $(K,L)$ be a pair of multicomplexes. Let $G\curvearrowright K$ be a simplicial action and let $H$ be a subgroup of $G$ which induces an action $H\curvearrowright(K,L)$. Assume that the orbits of $G$ in $L$ are induced by $H$. Then the rows of the following commutative diagram are both exact
[…]
where the vertical arrows are inclusions of cochains. In particular, if the inclusions $C^{\bullet}_{b}(K)^{G}\hookrightarrow C^{\bullet}_{b}(K)$ and $C^{\bullet}_{b}(L)^{H}\hookrightarrow C^{\bullet}_{b}(L)$ induce isomorphisms in cohomology, then $C^{\bullet}_{b}(K,L)^{G}\hookrightarrow C^{\bullet}_{b}(K,L)$ induces an isomorphism in cohomology.
2.5. The singular multicomplex
Let $X$ be a topological space. The singular multicomplex $\mathcal{K}(X)$ of $X$ is the multicomplex whose simplices are the singular simplices in $X$ with distinct vertices, up to affine parametrization [Gro82]. Therefore, the geometric realization $|\mathcal{K}(X)|$ of $\mathcal{K}(X)$ is a CW-complex with 0-skeleton corresponding to the set $X$ itself. Moreover, there is an obvious projection $S_{X}\colon|\mathcal{K}(X)|\rightarrow X$, which is a homotopy equivalence when $X$ is a CW-complex [FM23, Corollary 2.2]. The size of $\mathcal{K}(X)$ can be reduced without changing its homotopy type. We define a submulticomplex $\mathcal{L}(X)$ of $\mathcal{K}(X)$ as follows: the 0-skeleton of $\mathcal{L}(X)$ is the same as the one of $\mathcal{K}(X)$ (hence it corresponds to the set $X$); having already defined the $n$-skeleton $\mathcal{L}(X)^{n}$ of $\mathcal{L}(X)$, we define the $(n+1)$-skeleton by adding to $\mathcal{L}(X)^{n}$ one $(n+1)$-simplex for each homotopy class of $(n+1)$-simplices of $\mathcal{K}(X)$ whose facets are all contained in $\mathcal{L}(X)^{n}$. It turns out that the inclusion $|\mathcal{L}(X)|\hookrightarrow|\mathcal{K}(X)|$ is a homotopy equivalence [FM23, Theorem 3.23]. Even if the construction of $\mathcal{L}(X)$ depends on several choices, the submulticomplex $\mathcal{L}(X)$ is unique up to simplicial isomorphism [FM23, Theorem 3.23]. To the singular multicomplex we can associate an aspherical multicomplex $\mathcal{A}(X)$, which is defined as the quotient of $\mathcal{L}(X)$, where two simplices of $\mathcal{L}(X)$ are identified if and only if they share the same 1-skeleton. One can define a simplicial projection $\pi\colon\mathcal{L}(X)\rightarrow\mathcal{A}(X)$, which restricts to the identity on $\mathcal{L}(X)^{1}=\mathcal{A}(X)^{1}$. This quotient map induces an isomorphism on fundamental groups [FM23, Proposition 3.33] and kills higher homotopy [FM23, Theorem 3.31]. Therefore the topological realization $|\mathcal{A}(X)|$ of $\mathcal{A}(X)$ is a model for the classifying space of the fundamental group of $X$.
2.6. The group $\Pi(X,X)$ and its action on $\mathcal{A}(X)$
Let $X$ be a topological space. The group $\Pi(X,X)$, first introduced by Gromov in [Gro82], is defined as follows.
Definition 2.3.
Let $X_{0}$ be a subset of $X$. The set $\Omega(X,X_{0})$ consists of families of paths $\{\gamma_{x}\}_{x\in X_{0}}$ such that:
  • (1)
    for every $x\in X_{0}$, $\gamma_{x}\colon[0,1]\rightarrow X$ is a continuous path such that $\gamma_{x}(0)=x$ and $\gamma_{x}(1)\in X_{0}$;
  • (2)
    $\gamma_{x}$ is constant for all but finitely many $x\in X_{0}$;
  • (3)
    the map $X_{0}\rightarrow X_{0}$, $x\mapsto\gamma_{x}(1)$, is a bijection (with finite support).
The following concatenation of paths endows $\Omega(X,X_{0})$ with the structure of a semigroup: given two elements $\{\gamma_{x}\}_{x\in X_{0}}$ and $\{\gamma^{\prime}_{x}\}_{x\in X_{0}}$ of $\Omega(X,X_{0})$, their concatenation is $\{\gamma_{x}*\gamma^{\prime}_{\gamma_{x}(1)}\}_{x\in X_{0}}$, where * denotes the usual concatenation of paths. In order to obtain a group, we consider the set $\Pi(X,X_{0})$ of homotopy classes of elements of $\Omega(X,X_{0})$, where two elements $\{\gamma_{x}\}_{x\in X_{0}}$ and $\{\gamma^{\prime}_{x}\}_{x\in X_{0}}$ of $\Omega(X,X_{0})$ are homotopic if $\gamma_{x}$ is homotopic to $\gamma_{x}^{\prime}$ in $X$ relative to the endpoints, for every $x\in X_{0}$.
Elements of $\Pi(X,X)$ are usually denoted by listing the homotopically non-trivial paths in one of its representatives. If $V\subseteq U\subseteq X$, then the inclusion $(U,V)\hookrightarrow(X,X)$ induces a group homomorphism $\Pi(U,V)\rightarrow\Pi(X,X)$, whose image is denoted by $\Pi_{X}(U,V)$. If $X_{0}=\{x_{0}\}$, then $\Pi(X,X_{0})=\pi_{1}(X,x_{0})$ is just the fundamental group of $X$ at $x_{0}$. In general, we have an injective group homomorphism
$\bigoplus_{x\in X_{0}}\pi_{1}(X,x)\hookrightarrow\Pi(X,X_{0}),$
which is the kernel of the natural homomorphism from $\Pi(X,X_{0})$ to the group of permutations of $X_{0}$ with finite support [FM23, Proposition 6.5]. Since the latter is amenable, if every connected component of $X$ has amenable fundamental group, then $\Pi(X,X)$ is amenable. In general, we have the following.
Lemma 2.4 ([FM23, Lemma 6.6]).
Let $U$ be an amenable subset of $X$ and let $V\subseteq U$ be any subset. Then the subgroup $\Pi_{X}(U,V)\leq\Pi(X,X)$ is amenable.
The group $\Pi(X,X)$ acts on $\mathcal{A}(X)$ as follows. Let $g\in\Pi(X,X)$ and let $\{\gamma_{x}\}_{x\in X}$ be a representative of $g$. We define the action of $g$ on the 0-skeleton of $\mathcal{A}(X)$ as the permutation induced by $g$ on $\mathcal{A}(X)^{0}=X$. Let $e$ be an edge of $\mathcal{A}(X)$ with endpoints $v_{0},v_{1}\in\mathcal{A}(X)^{0}=X$. Recall that edges of $\mathcal{A}(X)$ (hence of $\mathcal{L}(X)$) correspond to homotopy classes (relative to endpoints) of paths in $X$. Let $\gamma_{e}\colon[0,1]\rightarrow X$ be a representative of $e$. We define $g\cdot e$ as the homotopy class (relative to the endpoints) of the path $\bar{\gamma}_{v_{0}}*\gamma_{e}*\gamma_{v_{1}}$. Since $\mathcal{A}(X)$ is aspherical, we can uniquely extend $g$ to a simplicial automorphisms of the whole $\mathcal{A}(X)$, which is simplicially homotopic to the identity [FM23, Theorem 5.3].
3. From pairs of spaces to pairs of multicomplexes
Let $(X,A)$ be a CW-pair. Being injective, the inclusion map $j\colon A\hookrightarrow X$ induces a simplicial embedding $j_{\mathcal{K}}\colon\mathcal{K}(A)\hookrightarrow\mathcal{K}(X)$ at the level of singular multicomplexes. We denote by $S_{X}\colon|\mathcal{K}(X)|\rightarrow X$ and $S_{A}\colon|\mathcal{K}(A)|\rightarrow A$ the corresponding projections. Of course we have that $S_{X}\circ|j_{\mathcal{K}}|=j\circ S_{A}$. We consider the multicomplexes $\mathcal{L}(A)$ and $\mathcal{L}(X)$, respectively, and we denote by $i_{A}\colon\mathcal{L}(A)\hookrightarrow\mathcal{K}(A)$ and $i_{X}\colon\mathcal{L}(X)\hookrightarrow\mathcal{K}(X)$ the corresponding simplicial inclusions. There is a simplicial map $j_{\mathcal{L}}\colon\mathcal{L}(A)\rightarrow\mathcal{L}(X)$ which sends each simplex $\Delta$ of $\mathcal{L}(A)$ to the unique simplex of $\mathcal{K}(X)$ which is homotopic to $\Delta$ relative to the 0-skeleton [Cap, Proposition 4.1]. Moreover, we can construct $\mathcal{L}(X)$ and $\mathcal{L}(A)$ such that $j_{\mathcal{K}}\circ i_{A}=i_{X}\circ j_{\mathcal{L}}$.
Remark 3.1.
Although the multicomplexes $\mathcal{L}(A)$ and $\mathcal{L}(X)$ are unique up to simplicial isomorphism [FM23, Theorem 3.23], some choices are needed in order to obtain that $j_{\mathcal{K}}\circ i_{A}=i_{X}\circ j_{\mathcal{L}}$. Namely, once $\mathcal{L}(A)$ has been constructed, we need to construct $\mathcal{L}(X)$ in the following way: if the facets of a simplex are all contained in $A$ and if there is a simplex of $\mathcal{L}(A)$ in the same homotopy class, we choose this as a representative. In this paper, whenever we construct a pair of multicomplexes from a CW-pair, we always adhere to this setting.
The map $j_{\mathcal{L}}$ is not injective in general: to this end one needs to have more control over the homotopy of the pair $(X,A)$ [Cap, Proposition 4.2]. Let $\pi\colon\mathcal{L}(X)\rightarrow\mathcal{A}(X)$ denote the simplicial projection which identifies simplices sharing the same 1-skeleton. Of course, $j_{\mathcal{L}}$ factors to a well-defined simplicial map $j_{\mathcal{A}}\colon\mathcal{A}(A)\rightarrow\mathcal{A}(X)$. We denote by $\mathcal{A}_{X}(A)$ the image of $\mathcal{A}(A)$ in $\mathcal{A}(X)$ via $j_{\mathcal{A}}$, so that the pair of multicomplexes $(\mathcal{A}(X),\mathcal{A}_{X}(A))$ is well-defined. We denote by $q\colon\mathcal{A}(A)\rightarrow\mathcal{A}_{X}(A)$ the surjective map induced by $j_{\mathcal{A}}$. If $A$ is $\pi_{1}$-injective in $X$, then $j_{\mathcal{A}}$ is a simplicial embedding and $\mathcal{A}_{X}(A)$ is simplicially isomorphic to $\mathcal{A}(A)$ [Kue15, Section 1.3] [Cap, Proposition 4.3]. In short, we have the following commutative diagram of simplicial maps
(3.1) $\mathcal{K}(A)$$\mathcal{L}(A)$$\mathcal{A}(A)$$\mathcal{A}_{X}(A)$$\mathcal{K}(X)$$\mathcal{L}(X)$$\mathcal{A}(X)$$\mathcal{A}(X).$$\scriptstyle{j_{\mathcal{K}}}$$\scriptstyle{i_{A}}$$\scriptstyle{j_{\mathcal{L}}}$$\scriptstyle{\pi}$$\scriptstyle{j_{\mathcal{A}}}$$\scriptstyle{q}$$\scriptstyle{i_{X}}$$\scriptstyle{\pi}$
Proposition 3.2 ([Cap, Theorem 7]).
Let $(X,A)$ be a CW-pair such that the kernel of the morphism $\pi_{1}(A\hookrightarrow X,x)$ is amenable for every $x\in A$. Then, for every $n\in\mathbb{N}$, there is an isomorphism of vector spaces
$\Psi^{n}\colon H^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))\rightarrow H^{n}_{b}(X,A).$
The previous result is achieved in [Cap] by using the machinery of mapping cones developed in [Par03, Löh08]. The map $\Psi^{n}$ is induced by the maps in (3.1), hence it is a priori not canonical, since our construction depends on the choices described in Remark 3.1. We refer to the next pages for the details of the construction, which can be also found in [Cap, Section 8].
3.1. Mapping cones
Let $f\colon L\rightarrow K$ be a simplicial map between multicomplexes. The mapping cone complex of $f$ is $(C^{\bullet}_{b}(f\colon L\rightarrow K),\bar{\delta}^{\bullet})$, where
$C^{n}_{b}(f\colon L\rightarrow K)=C^{n}_{b}(K)\oplus C^{n-1}_{b}(L),\quad\bar{\delta}^{n}(u,v)=(\delta^{n}(u),-f^{n}(u)-\delta^{n-1}(v)).$
The proof of the following is analogous to the one of [Par03, Theorem 3.19].
Lemma 3.3 ([Cap, Lemma 7.4]).
Let $(K,L)$ be a pair of multicomplexes and let $i\colon L\hookrightarrow K$ denote the inclusion map. The chain map $\beta^{\bullet}\colon C^{\bullet}_{b}(K,L)\rightarrow C^{\bullet}_{b}(i\colon L\hookrightarrow K)$, defined by $\beta^{n}(u)=(u,0)$, induces an isomorphism in cohomology in every degree.
The following lemma is a consequence of the naturality of the short exact sequence of chain complexes $0\rightarrow C^{\bullet-1}_{b}(L)\rightarrow C^{\bullet}_{b}(f\colon L\rightarrow K)\rightarrow C^{\bullet}_{b}(K)\rightarrow 0$.
Lemma 3.4.
For every commutative diagram of simplicial maps
$L{L^{\prime}}$$K{K^{\prime}}$$\scriptstyle{f}$$\scriptstyle{\xi_{L}}$$\scriptstyle{f^{\prime}}$$\scriptstyle{\xi_{K}}$
there is a chain map $\xi^{\bullet}\colon C^{\bullet}_{b}(f^{\prime}\colon L^{\prime}\rightarrow K^{\prime})\rightarrow C^{\bullet}_{b}(f\colon L\rightarrow K)$, defined by $\xi^{n}(u,v)=(\xi_{K}^{n}(u),\xi_{L}^{n-1}(v))$. Moreover, if both $\xi_{K}$ and $\xi_{L}$ induce isomorphisms in bounded cohomology, then also $\xi^{\bullet}$ does.
3.2. Proof of Proposition 3.2
Let $(X,A)$ be a CW-pair and consider the commutative diagram of simplicial maps (3.1). We know that $i_{X}$ and $i_{A}$ induce isomorphisms in bounded cohomology [FM23, Corollary 4.9]. The same holds for the map $\pi$ [FM23, Theorem 4.23]. Assume that the kernel of the morphism $\pi_{1}(A\hookrightarrow X,x)$ is amenable for every $x\in A$. This last assumption is crucial to show that also the map $q$ induces an isomorphism in bounded cohomology [Cap, Lemma 8.3]. It follows that every horizontal map of diagram (3.1) induces an isomorphism in bounded cohomology.
In the following, we implicitly invoke the Five Lemma several times to pass from the absolute case to the relative one. We know that the projection $S\colon(|\mathcal{K}(X)|,|\mathcal{K}(A)|)\rightarrow(X,A)$ is a homotopy equivalence both on $|\mathcal{K}(X)|$ and $|\mathcal{K}(A)|$ [FM23, Corollary 2.2]. If follows that the chain map $S^{\bullet}\colon C^{\bullet}_{b}(X,A)\rightarrow C^{\bullet}_{b}(|\mathcal{K}(X)|,|\mathcal{K}(A)|)$ induces an isomorphism in bounded cohomology. Let $\varphi_{\bullet}\colon C_{\bullet}(\mathcal{K}(X))\rightarrow C_{\bullet}(|\mathcal{K}(X)|)$ denote the natural chain inclusion. By the Isometry Lemma [FM23, Theorem 2], $\varphi_{\bullet}$ induces an isomorphism in bounded cohomology. Hence the map $\varphi^{\bullet}\colon C^{\bullet}_{b}(|\mathcal{K}(X)|,|\mathcal{K}(A)|)\rightarrow C^{\bullet}_{b}(\mathcal{K}(X),\mathcal{K}(A))$ induced by $\varphi_{\bullet}$ also does. By Lemma 3.3, the chain map $\beta^{\bullet}\colon C^{\bullet}_{b}(\mathcal{K}(X),\mathcal{K}(A))\rightarrow C^{\bullet}_{b}(j_{\mathcal{K}}\colon\mathcal{K}(A)\hookrightarrow\mathcal{K}(X))$ induces an isomorphism in cohomology. Since every horizontal map of diagram (3.1) induces an isomorphism in bounded cohomology, we can apply Lemma 3.4 to every square of diagram (3.1). It follows that the chain maps
$i^{\bullet}\colon C^{\bullet}_{b}(j_{\mathcal{K}}\colon\mathcal{K}(A)\hookrightarrow\mathcal{K}(X))$ $\rightarrow C^{\bullet}_{b}(j_{\mathcal{L}}\colon\mathcal{L}(A)\rightarrow\mathcal{L}(X)),$
$\pi^{\bullet}\colon C^{\bullet}_{b}(j_{\mathcal{A}}\colon\mathcal{A}(A)\rightarrow\mathcal{A}(X))$ $\rightarrow C^{\bullet}_{b}(j_{\mathcal{L}}\colon\mathcal{L}(A)\rightarrow\mathcal{L}(X)),$
$q^{\bullet}\colon C^{\bullet}_{b}(\mathcal{A}_{X}(A)\hookrightarrow\mathcal{A}(X))$ $\rightarrow C^{\bullet}_{b}(j_{\mathcal{A}}\colon\mathcal{A}(A)\rightarrow\mathcal{A}(X))$
induce isomorphisms in cohomology. Since $\mathcal{A}_{X}(A)$ is a subcomplex of $\mathcal{A}(X)$, we can apply Lemma 3.3 to get a chain map $\beta^{\bullet}\colon C^{\bullet}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))\rightarrow C^{\bullet}_{b}(\mathcal{A}_{X}(A)\hookrightarrow\mathcal{A}(X))$, which induces an isomorphism in cohomology.
In conclusion, both the maps $\zeta^{\bullet}=i^{\bullet}\circ\beta^{\bullet}\circ\varphi^{\bullet}\circ S^{\bullet}$ and $\vartheta^{\bullet}=\pi^{\bullet}\circ q^{\bullet}\circ\beta^{\bullet}$ induce isomorphisms in cohomology. For every $n\in\mathbb{N}$, the map $\Psi^{n}$ is then defined as the composition $H^{n}(\zeta^{\bullet})^{-1}\circ H^{n}(\vartheta^{\bullet})$.
4. Proof of Theorem 1
Let $(X,A)$ be a triangulable pair such that the kernel of the morphism $\pi_{1}(A\hookrightarrow X,x)$ is amenable for every $x\in A$. Let $(T,S)$ be a pair of simplicial complexes such that $(X,A)=(|T|,|S|)$.
There is a natural way to construct a copy of $T$ inside $\mathcal{A}(X)$ [FM23, Section 6.2]. Inside the singular multicomplex $\mathcal{K}(X)$ we can find a submulticomplex $\mathcal{K}_{T}(X)\cong T$ whose simplices correspond to equivalence classes of affine parametrizations of simplices of $T$. Moreover, we can assume that $\mathcal{K}_{T}(X)\subseteq\mathcal{L}(X)$, by choosing simplices of $\mathcal{K}_{T}(X)$ as representatives in their homotopy class. Since $T$ is a simplicial complex, every simplex in $\mathcal{K}_{T}(X)$ is uniquely determined by its 0-skeleton, hence the quotient map $\pi\colon\mathcal{L}(X)\rightarrow\mathcal{A}(X)$ is injective on $\mathcal{K}_{T}(X)$. Therefore we can construct a copy of $T$ inside $\mathcal{A}(X)$. The same construction applies to $S$, so that $\mathcal{K}_{S}(A)\subseteq\mathcal{L}(A)$. As before, both $\pi\colon\mathcal{L}(A)\rightarrow\mathcal{A}(A)$ and $q\colon\mathcal{A}(A)\rightarrow\mathcal{A}_{X}(A)$ are injective on $\mathcal{K}_{S}(A)$. It follows that $S$ sits inside $\mathcal{A}_{X}(A)\cap T$. Moreover, the choices which ensure that $\mathcal{K}_{T}(X)\subseteq\mathcal{L}(X)$ and $\mathcal{K}_{S}(A)\subseteq\mathcal{L}(A)$ are clearly compatible with the ones described in Remark 3.1, so that both the squares of the following diagram are commutative
(4.1) $\mathcal{K}_{S}(A)$$\mathcal{L}(A)$$\mathcal{K}(A)$$\mathcal{K}_{T}(X)$$\mathcal{L}(X)$$\mathcal{K}(X).$$\scriptstyle{j_{\mathcal{L}}}$$\scriptstyle{i_{A}}$$\scriptstyle{j_{\mathcal{K}}}$$\scriptstyle{i_{A}}$
We can therefore invoke Proposition 3.2 to get an isomorphism of vector spaces $\Psi^{n}\colon H^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))\rightarrow H^{n}_{b}(X,A)$. To this end, it is crucial that the kernel of the map $\pi_{1}(A\hookrightarrow X)$ is amenable. This is the only step in our proof in which we use this assumption.
Lemma 4.1.
Let $z\in C^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))$ be a bounded cocycle which vanishes on $C_{n}(T)\subseteq C_{n}(\mathcal{A}(X))$. Then $\mathrm{comp}^{n}\circ\Psi^{n}([z])=0$ in $H^{n}(X,A)$.
Proof.
We treat simplicial complexes as specific cases of multicomplexes, using the cohomology theory defined in Section 2.3. The inclusion of pairs $(T,S)\subseteq(\mathcal{A}(X),\mathcal{A}_{X}(A))$ induces a restriction map $r^{n}_{\mathcal{A}}\colon C^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))\rightarrow C^{n}_{b}(T,S)$. Similarly, by Lemma 3.3, the leftmost square of diagram (4.1) induces a restriction map $r^{n}_{\mathcal{L}}\colon C^{n}_{b}(\mathcal{L}(A)\rightarrow\mathcal{L}(X))\rightarrow C^{n}_{b}(S\hookrightarrow T)$ on mapping cones. We consider the following diagram
$C^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))$$C^{n}_{b}(T,S)$$C^{n}(T,S)$$C^{n}(T,S)$$C^{n}_{b}(\mathcal{L}(A)\rightarrow\mathcal{L}(X))$$C^{n}_{b}(S\hookrightarrow T)$$C^{n}(S\hookrightarrow T)$$C^{n}_{b}(X,A)$$C^{n}(X,A).$$\scriptstyle{r^{n}_{\mathcal{A}}}$$\scriptstyle{\zeta^{n}}$$\scriptstyle{\beta^{n}}$$\scriptstyle{\beta^{n}}$$\scriptstyle{r^{n}_{\mathcal{L}}}$$\scriptstyle{\vartheta^{n}}$$\scriptstyle{\varphi^{n}}$
Here hooked arrows denote inclusion of chains, which induce comparison maps in cohomology. The maps $\zeta^{n}$ and $\vartheta^{n}$, introduced in Section 3, induce isomorphisms in cohomology whose composition defines $\Psi^{n}$. The maps $\beta^{n}$, defined in Lemma 3.3, and the map $\varphi^{n}$, induced by the chain inclusion $C_{n}(T)\rightarrow C_{n}(|T|)=C_{n}(X)$, also induce isomorphisms in cohomology, by Lemma 3.3 and [FM23, Theorem 1.11] respectively. Since $z$ vanishes on $C_{n}(T)$, we have that $r^{n}_{\mathcal{A}}(z)=0$. Since the diagram is commutative and every vertical arrow induces isomorphism in cohomology, the statement easily follows. ∎
Let $\mathcal{U}=\{U_{i}\,|\,i\in I\}$ be an amenable open cover of $X$ by path-connected subsets satisfying conditions (RC1) and (RC2). We denote by $V$ the set of vertices of $T$. For every vertex $v\in V$, the closed star of $v$ in $T$ is the subcomplex of $T$ containing all the simplices containing $v$. By suitably subdividing $T$, we can assume that for every $v\in V$ there exists $i(v)\in I$ such that the closed star of $v$ in $T$ is contained in $U_{i(v)}$ [Mun84, Theorem 16.4]. Of course the choice of $i(v)$ may not be unique. For every $i\in I$, we set $V_{i}=\{v\in V(T)\,|\;i(v)=i\}$. We consider the group
$G=\bigoplus_{i\in I}\Pi_{X}(U_{i},V_{i}).$
Since $U_{i}$ is amenable in $X$, then $\Pi_{X}(U_{i},V_{i})$ is an amenable group by Lemma 2.4. The direct sum of amenable groups is amenable, hence also $G$ is amenable. Moreover, since $V_{i}\cap V_{j}=\emptyset$ for every $i\neq j$, elements of $\Pi_{X}(U_{i},V_{i})$ commute with elements of $\Pi_{X}(U_{j},V_{j})$. It follows that there is a well-defined map from $G$ to $\Pi(X,X)$. This map is injective because the intersection $\Pi_{X}(U_{i},V_{i})\cap\left\langle\Pi_{X}(U_{j},V_{j})\colon j\neq i\right\rangle$ is trivial for every $i\in I$. In fact, non-trivial elements of the former group are represented by paths with endpoints in $V_{i}$, while non-trivial elements of the latter by paths with endpoints in $\bigcup_{j\neq i}V_{i}$, which is disjoint from $V_{i}$. Therefore $G$ can be identified with a subgroup of $\Pi(X,X)$ and acts on $\mathcal{A}(X)$ accordingly. Notice that this action does not preserve the submulticomplex $\mathcal{A}_{X}(A)$. The group
$H=\bigoplus_{i\in I}\Pi_{X}(U_{i}\cap A,V_{i}\cap A)$
identifies a subgroup of $G$ and induces an action on the pair $(\mathcal{A}(X),\mathcal{A}_{X}(A))$. Being a subgroup of an amenable group, $H$ is amenable. The following lemma exploits the regularity conditions (RC1) and (RC2) satisfied by the open cover.
Lemma 4.2.
The action $G\curvearrowright\mathcal{A}(X)$ has orbits in $\mathcal{A}_{X}(A)$ induced by $H$.
Proof.
Let $\sigma=(\Delta,(x_{0},\dots,x_{n}))$ be an algebraic $n$-simplex of $\mathcal{A}_{X}(A)$ and let $g\in G$ be such that $g\cdot\sigma$ is an algebraic $n$-simplex of $\mathcal{A}_{X}(A)$. We need to show that there exists an element $h\in H$ such that $h\cdot\sigma=g\cdot\sigma$. Let $\{v_{0},\dots,v_{k}\}$ be the vertices of $\Delta$, where $k\leq n$, and let $\{\gamma_{x}\}_{x\in X}$ be any representative of $g$. Since $v_{j}$ and $g\cdot v_{j}$ are both vertices of $\mathcal{A}_{X}(A)$ (hence points of $A$), by the definition of the action $G\curvearrowright\mathcal{A}(X)$, we have that $\gamma_{v_{j}}$ has both endpoints in $A$. Moreover, by the definition of $G$, we have that $\gamma_{v_{j}}$ is supported in some element of $\mathcal{U}$. Hence there are $i_{0},\dots,i_{k}\in I$ such that $\gamma_{v_{j}}\colon[0,1]\rightarrow U_{i_{j}}$ has both endpoints in $U_{i_{j}}\cap A$. Since the open cover $\mathcal{U}$ satisfies (RC1) and (RC2), by Lemma 2.1 there exist paths $\lambda_{v_{j}}\colon[0,1]\rightarrow U_{i_{j}}\cap A$, $j\in\{0,\dots,k\}$, such that $\lambda_{v_{j}}$ is homotopic to $\gamma_{v_{j}}$ in $X$ relative to the endpoints. Therefore, if we set $h_{j}=\{\lambda_{j}\}$ if $\lambda_{j}$ is a loop, and $h_{j}=\{\lambda_{j},\bar{\lambda}_{j}\}$ otherwise, then $h=\oplus_{j=0}^{k}h_{j}$ defines an element of $H$. It is easy to check that $h\cdot e=g\cdot e$, for every edge $e$ of $\Delta$, hence $g\cdot\Delta^{1}=h\cdot\Delta^{1}$, where $\Delta^{1}$ denotes the 1-skeleton of $\Delta$. Since $\mathcal{A}(X)$ is an aspherical multicomplex, this is indeed sufficient to conclude [FM23, Proposition 3.30]. ∎
Let $n\geq\mathrm{mult}(\mathcal{U})$ and let $\alpha\in H^{n}_{b}(X,A)$. Since $\Psi^{n}$ is an isomorphism, we can take a bounded coclass $\beta\in H^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))$ such that $\Psi^{n}(\beta)=\alpha$. Let now $z\in C^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))$ be an alternating cocycle representing $\beta$. The actions $G\curvearrowright\mathcal{A}(X)$ and $H\curvearrowright\mathcal{A}_{X}(A)$ are both by amenable groups and by automorphisms which are simplicially homotopic to the identity [FM23, Theorem 5.3] [Cap, Proposition 6.1]. It follows from these two facts that the inclusion maps $C^{\bullet}_{b}(\mathcal{A}(X))^{G}\hookrightarrow C^{\bullet}_{b}(\mathcal{A}(X))$ and $C^{\bullet}_{b}(\mathcal{A}_{X}(A))^{H}\hookrightarrow C^{\bullet}_{b}(\mathcal{A}_{X}(A))$ induce isomorphisms in cohomology [FM23, Theorem 4.21]. Moreover, since the action $G\curvearrowright\mathcal{A}(X)$ has orbits in $\mathcal{A}_{X}(A)$ induced by $H$, by Lemma 2.2 we can assume that $z$ is alternating and $G$-invariant. By Lemma 4.1, it suffices to show that $z$ vanishes on every algebraic $n$-simplex of $C_{n}(T)\subseteq C_{n}(\mathcal{A}(X))$.
Let $(\Delta,(x_{0},\dots,x_{n}))$ be an algebraic $n$-simplex in $C_{n}(T)\subseteq C_{n}(\mathcal{A}(X))$. If $x_{h}=x_{k}$, for some $h\neq k$, then $z(\Delta,(x_{0},\dots,x_{n}))=0$, since $z$ is alternating. We assume therefore that the points $x_{0},\dots,x_{n}$ are pairwise distinct. By construction, the cover of $X$ by the closed star of vertices of $T$ is a refinement of $\mathcal{U}$. Since $\Delta$ is an $n$-simplex of $T$ and since $n\geq\mathrm{mult}(\mathcal{U})$, there are at least two vertices of $\Delta$ belonging to the same $V_{i}$. In other words, there exist $h\neq k\in\{0,\dots,n\}$ such that $i(x_{h})=i(x_{k})$. Let $e$ denote the edge of $\Delta$ joining $x_{h}$ with $x_{k}$. By assumption, the closed stars of $x_{k}$ and $x_{h}$ are both contained in $U_{i}$, therefore the edge $e$ of $\mathcal{A}(X)$ (which is also an edge of $\mathcal{L}(X)^{1}=\mathcal{A}(X)^{1}$, hence of $\mathcal{K}(X)$) projects via $S_{X}\colon|\mathcal{K}(X)|\rightarrow X$ to a path $\gamma\colon[0,1]\rightarrow U_{i}$ with endpoints $x_{h}$ and $x_{k}$. As a consequence, if we consider $g=\{\gamma,\bar{\gamma}\}\in\Pi_{X}(U_{i},V_{i})<G$, it is easy to check that $g\cdot\Delta=\Delta$, $g\cdot x_{h}=x_{k}$, $g\cdot x_{k}=x_{h}$ and $g\cdot x_{j}=x_{j}$, for every $j\neq h,k$. Since $z$ is $G$-invariant, we obtain $z(\Delta,(x_{0},\dots,x_{n}))=z(g\cdot(\Delta,(x_{0},\dots,x_{n})))$, while, since $z$ is alternating, we have $z(\Delta,(x_{0},\dots,x_{n}))=-z(g\cdot(\Delta,(x_{0},\dots,x_{n})))$. Therefore $z(s,(x_{0},\dots,x_{n}))=0$ and this concludes the proof.
Remark 4.3.
In our proof of Theorem 1 we work with a pair of simplicial complexes $(T,S)$ for two main technical reasons. The first is to exploit the barycentric techniques to relate stars of $T$ with the open cover $\mathcal{U}$. The second is to build a copy of $(X,A)$ inside $(\mathcal{A}(X),\mathcal{A}_{X}(A))$, on which $G$ acts.
Remark 4.4.
We discuss the role of conditions (RC1) and (RC2) in our context. In fact, in the light of Lemma 4.2, the only reason to introduce these hypothesis is that, under these regularity assumptions, one can show that the action of $G$ on $\mathcal{A}(X)$ has orbits in $\mathcal{A}_{X}(A)$ induced by $H$ (Lemma 4.2). This in turns allows to consider $G$-invariant cochains in the relative setting (Lemma 2.2). In fact, since the action of $G$ on $\mathcal{A}(X)$ does not preserve $\mathcal{A}_{X}(A)$, some care is needed when passing to invariant cochains. Assume, on the contrary, that one could work with $G$-invariant chains regardless of any regularity condition of the action of $G$ on $\mathcal{A}_{X}(A)$. This assumptions seem implicit in Kuessner’s work on relative bounded cohomology via multicomplexes (see for example [Kue15, Corollary 2]). In this case, using the same argument above, one could prove that the comparison map $\mathrm{comp}^{n}\colon H^{n}_{b}(X,A)\rightarrow H^{n}(X,A)$ vanishes for every $n\geq\mathrm{mult}(\mathcal{U})$, where $\mathcal{U}$ is any amenable open cover of $X$. This is obviously false: if $(X,A)=(M,\partial M)$ for some smooth $m$-manifold $M$ with non-empty boundary, then we know that $M$ is homotopy equivalent a subcomplex of dimension at most m1m-1, therefore it admits a contractible (hence amenable) open cover of cardinality at most $m$ [CLM22, Remark 2.8]; by the standard duality between bounded cohomology and simplicial volume [Fri17, Proposition 7.10], it would follow that $\|M,\partial M\|=0$. This gives a contradiction, for example, if the interior of $M$ admits a complete finite-volume hyperbolic metric [Gro82].
5. Proof of Theorem 2
Let $(X,A)$ be a triangulable pair such that the kernel of the morphism $\pi_{1}(A\hookrightarrow X,x)$ is amenable for every $x\in A$. Let $\mathcal{U}$ be an amenable open cover of $X$ by path-connected subsets satisfying (RC1) and (RC2). We keep the notation from the previous section.
5.1. Proof of (1)
We assume that $\mathcal{U}$ is weakly-convex on $A$. Using the fact that $\mathcal{U}$ is locally-finite, up to subdividing $T$, we may suppose that $V_{i}\cap A\neq\emptyset$ for every $i\in I$ such that $U_{i}\cap A\neq\emptyset$. Let $n\geq\mathrm{mult}_{A}(\mathcal{U})$. We need to show that the comparison map $\mathrm{comp}^{n}\colon H^{n}_{b}(X,A)\rightarrow H^{n}(X,A)$ vanishes. Let $\alpha\in H^{n}_{b}(X,A)$. Recall that we have the isomorphism $\Psi^{n}$, hence we can take a bounded coclass $\beta\in H^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))$ such that $\Psi^{n}(\beta)=\alpha$. Let $z\in C^{n}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))$ be a cocycle representing $\beta$. As in in the previous section, we can assume that $z$ is alternating and $G$-invariant. By Lemma 4.1, we need to show that $z$ vanishes on every algebraic $n$-simplex of $C_{n}(T)\subseteq C_{n}(\mathcal{A}(X))$.
Let $(\Delta,(x_{0},\dots,x_{n}))$ be an algebraic $n$-simplex in $C_{n}(T)$. Since $z$ is alternating, we can assume that the points $x_{0},\dots,x_{n}$ are pairwise-distinct. We set $i_{j}=i(x_{j})$ for every $j\in\{0,\dots,n\}$. If there exist $h\neq k\in\{0,\dots,n\}$ such that $i_{h}=i_{k}$, then we can argue as in the previous section to show that $z(\Delta,(x_{0},\dots,x_{n}))=0$. We assume therefore that $i_{0},\dots i_{n}$ are pairwise-distinct.
Since $n\geq\mathrm{mult}_{A}(\mathcal{U})$ and since by construction $\bigcap_{j=0}^{n}U_{i_{j}}\neq\emptyset$, we deduce that $U_{i_{0}}\cap\dots\cap U_{i_{n}}\cap A$ is nonempty. Hence, since $\mathcal{U}$ is weakly-convex on $A$, the path-connected component of $U_{i_{0}}\cap\dots\cap U_{i_{n}}$ containing $\Delta$ intersects $A$. Therefore, if $x$ denotes a barycenter of $\Delta$, there exists $x^{\prime}\in U_{i_{0}}\cap\dots\cap U_{i_{n}}\cap A$ and a path $\lambda\colon[0,1]\rightarrow U_{i_{0}}\cap\dots\cap U_{i_{n}}$ from $x^{\prime}$ to $x$. Let $e_{hk}$ denote the oriented 1-simplex of $\Delta$ from $x_{h}$ to $x_{k}$. We denote by $\delta_{j}\colon[0,1]\rightarrow U_{i_{0}}\cap\dots\cap U_{i_{n}}$ a path from $x_{j}$ to $x$ supported in $\Delta$ such that $\delta_{h}*\bar{\delta}_{k}$ is homotopic in $X$ relative to the endpoints to a parametrization of $e_{hk}$. Since $V_{i_{j}}\cap A\neq\emptyset$, for every $j\in\{0,\dots,n\}$ there exists $x_{j}^{\prime}\in V_{i_{j}}\cap A$. Moreover, by (RC1), $U_{i_{j}}\cap A$ is path-connected, hence we can find continuous paths $\xi_{j}\colon[0,1]\rightarrow U_{i_{j}}\cap A$ from $x_{j}^{\prime}$ to $x^{\prime}$. For $0\leq h<k\leq n$, we then set $\xi_{hk}=\xi_{h}*\bar{\xi}_{k}\colon[0,1]\rightarrow A$. Since $x_{0}^{\prime},\dots,x_{n}^{\prime}$ are pairwise-distinct, by definition of $\mathcal{A}_{X}(A)$, there exists a unique oriented 1-simplex $e_{hk}^{\prime}$ of $\mathcal{A}_{X}(A)$ whose projection is a path in $A$ which is homotopic to $\xi_{hk}$ in $X$ relative to the endpoints.
Lemma 5.1.
There exists a simplex $\Delta^{\prime}$ of $\mathcal{A}_{X}(A)$ whose 1-skeleton is given by the union of the $e_{hk}^{\prime}$.
Proof.
Notice that the $e_{hk}^{\prime}$ can be considered to be 1-simplices of $\mathcal{A}(A)$. Of course, the loop $\xi_{jh}*\xi_{hk}*\bar{\xi}_{jk}$ is null-homotopic in $A$ (hence in $X$). By [FM23, Proposition 3.33] the concatenation of oriented simplices $e_{jh}^{\prime}*e_{hk}^{\prime}*\bar{e}_{jk}^{\prime}$ is null-homotopic in $|\mathcal{A}(A)|$. By [FM23, Proposition 3.30], it follows that there exists a unique $n$-simplex $\Delta^{\prime\prime}$ of $\mathcal{A}(A)$ whose 1-skeleton is the union of the $e_{hk}^{\prime}$. In conclusion, $\Delta^{\prime}$ denotes just the image of $\Delta^{\prime\prime}$ in $\mathcal{A}_{X}(A)$. ∎
For every $j\in\{0,\dots,n\}$ we set $\gamma_{j}=\xi_{j}*\lambda*\bar{\delta}_{j}$. By construction $\gamma_{j}$ joins $x_{j}^{\prime}$ with $x_{j}$ and is supported on $U_{i_{j}}$. Therefore, if we set $g_{j}=\{\gamma_{j}\}$ if $x_{j}=x_{j}^{\prime}$, and $g_{j}=\{\gamma_{j},\bar{\gamma}_{j}\}$ if $x_{j}\neq x_{j}^{\prime}$, then $g=\oplus_{j=0}^{n}g_{j}$ defines an element of $G$. It is clear that $g\cdot e_{hk}^{\prime}=e_{hk}$. Since simplices of $\mathcal{A}(X)$ are determined by their 1-skeleton [FM23, Proposition 3.30], we obtain $g\cdot\Delta^{\prime}=\Delta$. Hence we have that $z(\Delta,(x_{0},\dots,x_{n}))=z(\Delta^{\prime},(x_{0}^{\prime},\dots,x_{n}^{\prime}))$, since $z$ is $G$-invariant, and $z(\Delta^{\prime},(x_{0}^{\prime},\dots,x_{n}^{\prime}))=0$, since $z$ vanishes on simplices of $\mathcal{A}_{X}(A)$. This concludes the proof of (1).
5.2. Proof of (2)
Assume that $\mathcal{U}$ is convex. We denote by $N(\mathcal{U})$ the nerve of $\mathcal{U}$ and by $N(\mathcal{U}_{A})$ the nerve of the open cover of $A$ induced by $\mathcal{U}$. This is a well-defined subcomplex of $N(\mathcal{U})$ under our assumptions. Given a partition of unity subordinated to $\mathcal{U}$, one can construct a nerve map $\nu\colon X\rightarrow|N(\mathcal{U})|$, which is unique up to homotopy and which induces a well-defined map of pairs $\nu\colon(X,A)\rightarrow(|N(\mathcal{U})|,|N(\mathcal{U}_{A})|)$. By assumption, we know that $\mathcal{U}$ is locally-finite. Therefore, up to taking further subdivisions of $T$, we may suppose that $V_{i}\neq\emptyset$, for every $i\in I$, and $V_{i}\cap A\neq\emptyset$ for every $i\in I$ such that $U_{i}\cap A\neq\emptyset$. We need to show that there exists a map $\Theta^{n}\colon H^{n}_{b}(X,A)\rightarrow H^{n}(N(\mathcal{U}),N(\mathcal{U}_{A}))$ such that the following diagram is commutative:
(5.1) $H^{n}_{b}(X,A)$$H^{n}(X,A)$$H^{n}(N(\mathcal{U}),N(\mathcal{U}_{A}))$$H^{n}(|N(\mathcal{U})|,|N(\mathcal{U}_{A})|).$$\scriptstyle{\mathrm{comp}^{n}}$$\scriptstyle{\Theta^{n}}\scriptstyle{\cong}$$\scriptstyle{H^{n}(\nu)}$
Recall from the previous sections that the relative bounded cohomology of $(\mathcal{A}(X),\mathcal{A}_{X}(A))$ may be computed by the complex $C^{\bullet}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))^{G}_{\textup{alt}}$ of $G$-invariant alternating cochains which vanish on simplices supported on $\mathcal{A}_{X}(A)$ (see Lemma 2.2 and Lemma 4.2). Therefore, in order to define $\Theta^{n}$, we first construct chain maps $\Omega_{X}^{\bullet}$ and $\Omega_{A}^{\bullet}$ such that the following diagram is commutative
(5.2) $C^{\bullet}_{b}(\mathcal{A}(X))^{G}_{\textup{alt}}$$C^{\bullet}_{b}(\mathcal{A}_{X}(A))^{H}_{\textup{alt}}$$C^{\bullet}_{b}(N(\mathcal{U}))$$C^{\bullet}_{b}(N(\mathcal{U}_{A})),$$\scriptstyle{\Omega_{X}^{\bullet}}$$\scriptstyle{\Omega_{A}^{\bullet}}$
where the horizontal arrows are restriction maps. The map $\Omega_{X}^{n}$ can be constructed verbatim as in [FM23, Section 6.4]. The construction of $\Omega_{A}^{n}$ goes as follows. We identify the set of vertices of $N(\mathcal{U}_{A})$ with those indices $i\in I$ such that $U_{i}\cap A\neq\emptyset$. Let $z\in C^{n}_{b}(\mathcal{A}_{X}(A))^{H}_{\textup{alt}}$ and let $(i_{0},\dots,i_{n})\in N_{A}(\mathcal{U})$. If there exists $h\neq k\in\{0,\dots,n\}$ such that $i_{h}=i_{k}$, we set $\Omega_{A}^{n}(z)(i_{0},\dots,i_{n})=0$. Otherwise, by definition of the nerve $N(\mathcal{U}_{A})$, we have that $U_{i_{0}}\cap\dots\cap U_{i_{n}}\cap A\neq\emptyset$, hence we may choose a point $q\in U_{i_{0}}\cap\dots\cap U_{i_{n}}\cap A$. Moreover, since $V_{i_{j}}\cap A\neq\emptyset$ for every $j\in\{0,\dots,n\}$, we can pick a point $v_{i_{j}}\in V_{i_{j}}\cap A$. We know by (RC1) that $U_{i_{j}}\cap A$ is path-connected. Hence there exists a path $\alpha_{j}\colon[0,1]\rightarrow U_{i_{j}}\cap A$ from $v_{i_{j}}$ to $q$. For $0\leq h<k\leq n$, we set $\alpha_{hk}=\alpha_{h}*\bar{\alpha}_{k}$. Since $v_{i_{0}},\dots,v_{i_{n}}$ are pairwise distinct, by definition of $\mathcal{A}_{X}(A)$, there exists a unique oriented 1-simplex $e_{hk}$ of $\mathcal{A}_{X}(A)$ whose projection on $A$ is a path which is homotopic to $\alpha_{hk}$ in $X$ relative to the endpoints. Using the same argument of Lemma 5.1, there exists a unique $n$-simplex $\Delta$ of $\mathcal{A}_{X}(A)$ whose 1-skeleton is given by the union of the $e_{hk}$. We then set $\Omega^{n}_{A}(z)(i_{0},\dots,i_{n})=z(\Delta,(v_{i_{0}},\dots,v_{i_{n}}))$.
We need to show that $\Omega^{n}_{A}(z)$ is well-defined, i.e. different choices in the construction lead to the same value for $z(\Delta,(v_{i_{0}},\dots,v_{i_{n}}))$. Let $q^{\prime}\in U_{i_{0}}\cap\dots\cap U_{i_{n}}\cap A$, let $v_{i_{j}}^{\prime}\in V_{i_{j}}\cap A$, $j\in\{0,\dots,n\}$, and let $\alpha_{j}^{\prime}\colon[0,1]\rightarrow U_{i_{j}}\cap A$ be a path from $v_{i_{j}}^{\prime}$ to $q^{\prime}$. We set $\alpha_{hk}^{\prime}=\alpha_{h}^{\prime}*\bar{\alpha}_{k}^{\prime}$ and we take $e_{hk}^{\prime}$ to be the unique edge of $\mathcal{A}_{X}(A)$ whose projection is a path in $A$ which is homotopic to $\alpha_{hk}^{\prime}$ in $X$ relative to the endpoints. Finally, we take $\Delta^{\prime}$ to be the $n$-simplex of $\mathcal{A}_{X}(A)$ whose 1-skeleton is given by the $e_{hk}^{\prime}$. Since $U_{i_{0}}\cap\dots\cap U_{i_{n}}$ is path-connected (by convexity of $\mathcal{U}$), there exists a path $\beta\colon[0,1]\rightarrow\bigcap_{j=0}^{n}U_{i_{j}}$ from $q$ to $q^{\prime}$. By construction, the path $\gamma_{j}\coloneqq\alpha_{j}*\beta*\bar{\alpha}_{j}$ is supported on $U_{i_{j}}$ and joins $v_{i_{j}}$ with $v_{i_{j}}^{\prime}$, which are both points of $A$. By the regularity conditions (RC1) and (RC2), we deduce from Lemma 2.1 that, for every $j\in\{0,\dots,n\}$, there exists a path $\lambda_{j}\colon[0,1]\rightarrow U_{i_{j}}\cap A$ which is homotopic to $\gamma_{j}$ in $X$ relative to the endpoints. Therefore, if we set $h_{j}=\{\lambda_{j}\}$ if $\lambda_{j}$ is a loop, and $h_{j}=\{\lambda_{j},\bar{\lambda}_{j}\}$ otherwise, then $h=\oplus_{j=0}^{n}h_{j}$ defines an element of $H$. It is straightforward to check that $h\cdot e_{hk}^{\prime}=e_{hk}$ for every $h\neq k$, which implies that $h\cdot\Delta^{\prime}=\Delta$. Since $z$ is $H$-invariant, it follows that $z(\Delta,(v_{i_{0}},\dots,v_{i_{n}}))=z(\Delta^{\prime},(v_{i_{0}}^{\prime},\dots,v_{i_{n}}^{\prime}))$ i.e. that $\Omega_{A}^{n}$ is indeed well-defined.
It is easy to check that (5.2) is commutative and that $\Omega_{A}^{\bullet}$ defines a chain map. From (5.2), we get a chain map
$\Omega^{\bullet}\colon C^{\bullet}_{b}(\mathcal{A}(X),\mathcal{A}_{X}(A))^{G}_{\textup{alt}}\rightarrow C^{\bullet}(N(\mathcal{U}),N(\mathcal{U}_{A})),$
whose induced map in cohomology, precomposed with the inverse of $\Psi^{n}$, defines $\Theta^{n}$. The commutativity of (5.1) can be checked verbatim as in the absolute case (see [FM23, Section 6.4]). This concludes the proof of (2).